OpenStudy (anonymous):

x+2y=6 (1) x=2-2y (2) What is the solution of the system?

OpenStudy (anonymous):

is the ordered pair (2,-4) ??

OpenStudy (anonymous):

x+2y=6 (1) x=2-2y (2) Make 2y the subject in both 2y=6 - x(1) 2y = 2-x (2) Subtract [2] from [1] 0 = 4 (which we both know is nonesense) Therefore no real solutions as they are parallel lines (x values identical in both equations)

OpenStudy (anonymous):

oooh crap, well thanks lol

OpenStudy (anonymous):

i figured it out thank you :D

OpenStudy (anonymous):

niente!